# Actminer: Applying Causality Tracking and Increment Aligning for Graph-based Threat Hunting

### v1

@@ Abstract
To defend against Advanced Persistent Threats on the endpoint, threat hunting employs security knowledge such as cyber threat intelligence to continuously analyze system audit logs through retrospective scanning, querying, or pattern matching, aiming to uncover attack patterns/graphs that traditional detection methods (e.g., recognition for Point of Interest) fail to capture. However, existing threat hunting systems based on provenance graphs face challenges of high false negatives, high false positives, and low efficiency when confronted with diverse attack tactics and voluminous audit logs.
@@
To address these issues, we propose a system called Actminer, which constructs query graphs from descriptive relationships in cyber threat intelligence reports for precise threat hunting (i.e., graph alignment) on provenance graphs. First, we present a heuristic search strategy based on equivalent semantic transfer to reduce false negatives. Second, we establish a filtering mechanism based on causal relationships of attack behaviors to mitigate false positives. Finally, we design a tree structure to incrementally update the alignment results, significantly improving hunting efficiency. Evaluation on the DARPA Engagement dataset demonstrates that compared to the SOTA POIROT, Actminer reduces false positives by 39.1%, eliminates all false negatives, and effectively counters adversarial attacks.
Index Terms:
Threat Hunting, Advanced Persistent Threat, Attack Scenario Graph, Data Provenance.
I Introduction
@@
Advanced Persistent Threats (APTs) aim to infiltrate specific institutions to obtain critical asset information and sensitive data, posing immense threats and impacts. In order to detect and investigate APT attacks on hosts, data provenance [1] is widely used to analyze basic events (e.g., a sensitive file written by a malicious process) step-by-step. Nowadays, there are increasingly researchers have begun to employ provenance graphs in the field of APT attack detection on hosts.
@@
Existing provenance-based systems for detecting APTs fall into the following two main categories: rule-based [2, 3, 4, 5, 6, 7, 8, 9, 10], and learning-based [11, 12, 13, 14, 15, 16, 17]. The policies presented by rule-based systems are difficult to sustain in a constantly changing system environment, analysts must frequently update the rule base to adapt to new types of attacks, and the lag in defense results in frequent occurrences of false negatives. From the perspective of detection granularity, learning-based detection methods can be categorized into graph-level [14, 15, 11, 17], node/edge-level [12, 13, 16]. Graph-level detection typically involves learning the characteristics of benign graphs and use certain technique like clustering to discern abnormal ones. However, in the anomalous subgraph [17, 14, 15], not all nodes/edges are necessarily associated with the attack. Experts still need further analysis to pinpoint the malicious attack path accurately. In contrast, node/edge-level detection methods are able to obtain the point of interesting (POI), which makes it more direct and effective than graph-level ones in locating the attack candidate. However, the alerts that solely focus on nodes and edges do not fully reveal the panorama of the attack, and analysts still need to spend a significant amount of time evaluating whether the generated POIs are false positives [13, 12, 16]. According to CrowdStrike’s 2024 Global Threat Report [18], the time to compromise a host has gone from 84 minutes in 2022 to 62 minutes in 2023. This means that if an attack is not detected and responded to in a timely manner, the attacker will likely have a lateral movement [19] that will cause more hazards.
@@
Considering the shortcomings of the above detection systems, POIROT [2] has proposed a threat hunting approach based on graph alignment. It extracts the query graph from Cyber threat intelligence reports and then performs graph matching on the provenance graph to capture malicious behaviors. Due to the inherent interpretability of the query graph, its matching results can reflect comprehensive information about the attack, enabling analysts to respond accurately and promptly to the attack.
But POIROT still faces the following three limitations:
@@
- Temporal & Causality Missing (C2). How to detect causal relationships of attacks in dynamic scenarios to minimize erroneous alerts. Solely concentrating on a single behavior can overwhelm the hunting system with an influx of alerts, thereby hindering its ability to effectively identify and mitigate truly malicious activities (e.g., discerning the semantic differences in the access to sensitive files such as $/etc/passwd$ between normal processes and malicious processes requires consideration of causal relationships). This disregard may lead to more imprecise detection, rendering the system ineffective in countering sophisticated threats.
@@
In this work, we propose Actminer, a threat hunting system that combines causality tracking and incremental aligning to efficiently and accurately dig attack chains. To tackle C1, Actminer constructs a heuristic search strategy based on equivalent semantic transfer to counter phenomena such as attack camouflage, persistence, and evasion. We fuse the data information of inter-entity interactions through entities and their contextual semantics in order to achieve the accurate capture of malicious behaviors. To address C2, Actminer constructs a filtering mechanism based on the causal relationships of attack behaviors, and ignores unreasonable entity context relationships. Actminer employs the causal motivation behind attacks to guide threat hunting, ensuring the interpretability of hunting results and minimizing false positives. In other words, we provide a more accurate hunting result by excluding unreasonable (attack-irrelevant) paths based on the causal relationship through temporal sequences. To deal with C3, Actminer construct a tree structure to incrementally update the alignment results, thereby avoiding the significant overhead caused by rescanning multiple batches of redundant data.
We evaluate the effectiveness and efficiency of Actminer on the dataset provided by Darpa TC program [20, 21]. Our results reveal that Actminer surpasses existing provenance-based threat hunting system in terms of detection precision and recall. Moreover, Actminer can reduce the computational overhead and eliminate redundant searches. By deploying Actminer, security analysts are able to effectively analyze attack chains and formulate countermeasures, significantly alleviating the workload. In summary, the main contributions of our work are as follows:
  • Unlike traditional attack detection methods, we propose a provenance-based threat hunting system Actminer to accurately capture attack chains.
  • We introduce a heuristic search strategy based on equivalent semantic transfer and a filtering mechanism based on causal relationships of attack behaviors to ensure the precision and recall of Actminer.
@@
  • We comprehensively evaluate our system and SOTA POIROT [2] on the dataset from DARPA TC program. The results demonstrate the efficiency of Actminer in capturing the attack chain and highlight its resistance to adversarial attacks.
II Background Knowledge
II-A Provenance Graph
@@
Provenance graphs possess potent semantic expressiveness and contextual association capabilities, embodying the concrete manifestation of kernel audit logs. They model all system entities within the logs as nodes and the interactions between entities as edges, where both nodes and edges bear attribute information. The nodes within the provenance graph are categorized into subjects and objects based on the direction of data movement. Edges represent the causal relationships between system entities, such as read/write file operations, execute executable file operations, create/clone process operations, and so forth. By leveraging provenance graphs, security professionals can associate malicious entities with attack behaviors through causal analysis, unveiling the complete picture of an attack.
II-B CTI Report and Query Graph
@@
Cyber threat intelligence (CTI) reports [22, 23, 24] encompass comprehensive information related to cyber attacks and attackers, with a particular emphasis on capturing detailed attack procedures - the intricate sequences of steps and techniques employed in multi-stage attacks. These reports provide in-depth representations of attack scenarios, potential impacts on target hosts, as well as the complex chains of causally-linked events that characterize APTs. Security professionals leverage CTI reports to formulate more targeted defense rules for preventing and identifying malicious attack behaviors. In recent years, substantial research [11, 2, 25, 26, 27], has demonstrated the successful application of CTI reports in threat detection and threat hunting. In this paper, we construct a directed graph, termed the query graph, from the offensive and defensive knowledge[28](attack entities and their causal relationships) extracted from CTI reports. Similar to provenance graphs, query graphs are directed graphs with attribute information.
II-C Graph Alignment
@@
Graph alignment refers to the problem of detecting potential cyber intrusion behaviors by establishing an optimal subgraph mapping between a provenance graph ($G_{p}$) representing system activities across the entire system, and a query graph ($G_{q}$) representing attack pattern activities. The provenance graph $G_{p}$ = ($V_{p}$, $E_{p}$) consists of a node set $V_{p}$ representing system entities and events, and an edge set $E_{p}$. The query graph $G_{q}$ = ($V_{q}$, $E_{q}$) comprises a node set $V_{q}$ representing attack patterns and an edge set $E_{q}$. The goal of graph alignment is to find a subgraph $G_{m}$ in $G_{p}$ that maximizes the matching degree with $G_{q}$:
@@
Here, $M$ is the Matches function that calculates the matching degree between $G_{q}$ and a subgraph $G^{\prime}$ of $G_{p}$. By solving this optimization problem, the best mapping from the attack query graph to the activity graph is obtained, enabling the detection and tracking of cyber intrusion behaviors.
III Motivation
III-A Motivating Example
@@
As illustrated in Figure 1, this example includes two graphs: the top-left depicts the attack query graph manually extracted from a cyber threat intelligence (CTI) report, following the approach outlined in the POIROT.
The attack initiates by leveraging a malicious executable, malicious.exe, to obtain unauthorized system access. It then employs Registry modifications to establish persistence. Once a foothold is secured, the attacker can remotely issue commands and execute them on the system, executing an $art.bat_{2}$ file in the temporary folder and facilitating actions such as exfiltrating sensitive data to an external IP, exemplified by the transfer of /etc/passwd containing user account information.
On the right side is the provenance graph constructed from actual system logs capturing the observed execution behavior. Due to the fragmented nature of attack scenarios, coupled with the constraint of limiting the hop count to existing threat hunting approaches, can lead to imprecise or incomplete results during the threat hunting process. In this paper, we transform the threat-hunting problem into finding the attack query graph within the provenance graph.
@@
False Negative. Due to the complexity of real enterprise environments, semantic gaps exist between provenance graphs and attack query graphs. The manifestations of the same attack type may differ across systems, and attackers may utilize the same tools in diverse ways. For instance, entity names in the attack query graph may have varying representations in the underlying logs of different systems. In POIROT, regular expressions are employed to instantiate node names from the attack query graph for hunting searches in the provenance graph. However, if attackers modify their tactics, introducing technical variations, threat hunting systems struggle to detect different mutated attack samples (e.g., over 100 versions of the Carbanak malware were described in CTI reports). Furthermore, attackers can evade security detection through obfuscation, persistence, and evasion techniques. As illustrated, the attack query graph only describes data exfiltration over the network, whereas in the real environment, the attacker creates processes to read files, followed by another process reading those files, and finally transmitting them over the network. Overreliance on the simplistic approach of threat hunting based on a predetermined number of hops may inadvertently overlook malicious activities that align with the intrinsic characteristics of attacks, ultimately resulting in detection efforts failure. As shown in Figure 1, in a simulated scenario, as described in [2] Section 5, we set $C_{thr}$ to 3, but find that this limitation resulted in an incomplete capture of the attack chain. Consequently, it was unable to detect the art_bat file. Moreover, with such hop count restrictions, attackers aware of the imposed limits could potentially evade hunting more easily across different scenarios, resulting in potential harm. The manual adjustment of the hop limit according to different scenarios poses significant challenges. Similarly, the same problem exists in other path-based detection [29, 30, 31] efforts.
@@
False Positive. Within real organizations, extensive legitimate user operations exhibit similarities with attack behaviors in log data. If hunting rules are overly broad or incomplete, normal behaviors may be misclassified as malicious. For example, when a user downloads network files through a browser, the browser collects user data and transmits it to its cloud server, while the downloaded network files may be flagged as ”suspicious files” by the system, resembling malicious attack behaviors and triggering false alarms from the hunting system. Additionally, attackers may leverage tools/techniques to deceive hunting systems, also leading to false positives. As highlighted in Figure 1, a suspicious process Mal.exe exhibits two paths for reading/writing sensitive files. According to the attack query graph, sensitive file access should occur before network transmission, while $path_{1}$ occurs after transmission. Therefore, $path_{2}$ represents the attacker’s actual operations in the environment.
@@
High overhead and inefficiency. Government and enterprise organizations typically need to collect data simultaneously from thousands of machines, easily amassing petabyte-scale data volumes. This massive data not only imposes substantial storage overhead but also significantly reduces hunting efficiency. Traditional hunting methods require offline storage and continuous backscanning of system log data, resulting in immense computational overhead for each hunting operation. Referring to Figure 1, assume that all operations before node E23.txt at time 114 have already occurred. When a security analyst attempts to hunt for threats solely based on the data collected after this time, the incremental data segment alone cannot effectively support the reconstruction of the complete attack chain represented by its query graph. When examining the issue holistically, the newly acquired data lacks the necessary evidence to capture the earlier stages of the multi-step intrusion. Consequently, subsequent data collection would necessitate rescanning the previously available information, redundantly recomputing the provenance of data prior to a specific timeframe. These redundant computations across multiple hunting activities introduce an unsustainable overhead, hindering the system’s efficiency and scalability. And potentially allowing malicious activities to persist undetected for extended periods.
IV System Design

### v2

@@ Abstract
To defend against advanced persistent threats on the endpoint, threat hunting employs security knowledge, such as cyber threat intelligence (CTI) , to continuously analyze system audit logs through retrospective scanning, querying, or pattern matching, aiming to uncover attack patterns/graphs that traditional detection methods (e.g., recognition for point of interest) fail to capture. However, existing threat hunting systems based on provenance graphs face challenges of high false negatives (FNs), high false positives (FPs), and low efficiency when confronted with diverse attack tactics and voluminous audit logs. To address these issues, we propose a system called Actminer, which constructs query graphs from descriptive relationships in CTI reports for precise threat hunting (i.e., graph alignment) on provenance graphs. First, we present a heuristic search strategy based on equivalent semantic transfer to reduce FNs. Second, we establish a filtering mechanism based on causal relationships of attack behaviors to mitigate FPs. Finally, we design a tree structure to incrementally update the alignment results, significantly improving hunting efficiency. Evaluation on the DARPA Engagement dataset demonstrates that compared with the SOTA POIROT, Actminer reduces FPs by 39.1%, eliminates all FNs, and effectively counters adversarial attacks.
keywords:
[…]
1 Introduction
@@
Advanced persistent threats (APTs) aim to infiltrate specific institutions to obtain critical asset information and sensitive data, posing immense threats and impacts. Data provenance bates2015trustworthy is widely used to analyze basic events (e.g., a sensitive file written by a malicious process) in a step-by-step manner to detect and investigate APT attacks on hosts. Nowadays, researchers have increasingly begun to employ provenance graphs in the field of APT attack detection on hosts.
@@
Existing provenance-based systems for detecting APTs fall into the following two main categories: rule-based milajerdi2019poirot ; milajerdi2019holmes ; hossain2020combating ; xie2018pagoda ; bishop2005introduction ; kruegel2004intrusion ; xiong2020conan ; zhu2023aptshield ; hossain2017sleuth and learning-based altinisik2023provg ; zengy2022shadewatcher ; wang2022threatrace ; han2020unicorn ; manzoor2016fast ; rehman2024flash ; yang2023prographer systems. The policies presented by rule-based systems are difficult to sustain in a constantly changing system environment, analysts must frequently update the rule base to adapt to new types of attacks, and the lag in defense results in frequent occurrences of false negatives (FNs). From the perspective of detection granularity, learning-based detection methods can be categorized into graph-level han2020unicorn ; manzoor2016fast ; altinisik2023provg ; yang2023prographer and node/edge-level zengy2022shadewatcher ; wang2022threatrace ; rehman2024flash detection. Graph-level detection typically involves learning the characteristics of benign graphs and using certain techniques, such as clustering, to discern abnormal ones. However, in an anomalous subgraph yang2023prographer ; han2020unicorn ; manzoor2016fast , not all nodes/edges are necessarily associated with the attack. Thus, experts still need further analysis to accurately pinpoint the malicious attack path. In contrast, node/edge-level detection methods can obtain the point of interest (POI), making them more direct and effective than graph-level detection methods in locating the attack candidate. However, alerts that solely focus on nodes and edges do not fully reveal the panorama of the attack, and analysts still need to spend a significant amount of time evaluating whether the generated POIs are false positives (FPs). According to CrowdStrike’s 2024 Global Threat Report CrowdStrike2024 , the time to compromise a host has gone from 84 min in 2022 to 62 min in 2023. This means that if an attack is not detected and responded to in a timely manner, the attacker will likely perform lateral movement lateral-movement , which will cause more hazards.
@@
Considering the shortcomings of the above detection systems, POIROT milajerdi2019poirot has proposed a threat hunting approach based on graph alignment. This approach extracts the query graph from cyber threat intelligence (CTI) reports and then performs graph matching on the provenance graph to capture malicious behaviors. Due to the inherent interpretability of the query graph, its matching results can reflect comprehensive information about the attack, enabling analysts to respond accurately and promptly to the attack.
However, POIROT still faces the following three limitations:
@@
- Temporal and Causality Missing (C2). How to detect causal relationships of attacks in dynamic scenarios to minimize erroneous alerts. Solely concentrating on a single behavior can overwhelm the hunting system with an influx of alerts, thereby hindering its ability to effectively identify and mitigate truly malicious activities (e.g., discerning the semantic differences in the access to sensitive files, such as /etc/passwd/etc/passwd, between normal processes and malicious processes requires consideration of causal relationships). This disregard may lead to more imprecise detection, rendering the system ineffective in countering sophisticated threats.
@@
In this work, we propose Actminer, a threat hunting system that combines causality tracking and incremental aligning to efficiently and accurately dig attack chains. To tackle C1, Actminer constructs a heuristic search strategy based on equivalent semantic transfer (EST) to counter phenomena such as attack camouflage, persistence, and evasion. We fuse the data information of inter-entity interactions through entities and their contextual semantics to accurately capture malicious behaviors. To address C2, Actminer constructs a filtering mechanism based on the causal relationships of attack behaviors, and ignores unreasonable entity context relationships. Actminer employs the causal motivation behind attacks to guide threat hunting, ensuring the interpretability of hunting results and minimizing FPs. In other words, we provide a more accurate hunting result by excluding unreasonable (attack-irrelevant) paths based on the causal relationship through temporal sequences. To deal with C3, Actminer constructs a tree structure to incrementally update the alignment results, thereby avoiding the significant overhead caused by rescanning multiple batches of redundant data.
We evaluate the effectiveness and efficiency of Actminer on the dataset provided by DARPA-TC program DARPA3 ; darpa-tc . Our results reveal that Actminer surpasses existing provenance-based threat hunting systems in terms of detection precision and recall. Moreover, Actminer reduces the computational overhead and eliminates redundant searches. By deploying Actminer, security analysts can effectively analyze attack chains and formulate countermeasures, significantly alleviating the workload. In summary, the main contributions of our work are as follows:
  • Unlike traditional attack detection methods, we propose a provenance-based threat hunting system, i.e., Actminer, to accurately capture attack chains.
  • We introduce a heuristic search strategy based on EST and a filtering mechanism based on causal relationships of attack behaviors to ensure the precision and recall of Actminer.
@@
  • We comprehensively evaluate our system, POIROT milajerdi2019poirot , and MEGR-APT MEGR-APT on the dataset from the DARPA-TC program. The results demonstrate the efficiency of Actminer in capturing the attack chain, highlighting its resistance to adversarial attacks.
2 Background Knowledge
2.1 Advanced Persistent Threats and Threat Hunting
Advanced Persistent Threats (APTs) are sophisticated cyberattacks characterized by their covert nature, targeted focus, and emphasis on data extraction. APTs are meticulously planned and involve multiple phases aimed at infiltrating and maintaining persistence in the target system undetected. APT actors select tools tailored to the target environment and the objectives of their mission milajerdi2019holmes ; xiong2020conan .
The capabilities of APT actors have evolved over time as these groups expand their target scope, thereby adopting enhanced or new Tactics, Techniques, and Procedures (TTPs). These groups can develop tailored malware and data exfiltration methods. APT threat modeling is an extension of general threat modeling. It recognizes that attacks or campaigns are typically conducted through a series of phases. The Cyber Kill Chain lockheedmartinckc and MITRE’s ATT&CK framework MITREATTCK were developed based on extensive observations of APTs and related threat intelligence.
Threat hunting is a proactive cybersecurity approach that capitalizes on the distinct characteristics of APTs to detect and mitigate these threats. APTs follow a multi-phase attack lifecycle, as outlined in models like the Cyber Kill Chain and MITRE ATT&CK framework. Threat hunters analyze system logs, network traffic, and endpoint data to identify indicators of compromise (IOCs) or anomalous behaviors that match known APT TTPs. Recently, substantial research altinisik2023provg ; MEGR-APT ; milajerdi2019poirot ; mei2021ctscopy ; sun2023cyber has demonstrated that by analyzing the characteristics of TTPs intelligence associated with APT attacks, it is possible to proactively prevent such attacks from occurring.
2.2 Provenance Graph
@@
Provenance graphs possess potent semantic expressiveness and contextual association capabilities, embodying the concrete manifestation of kernel audit logs. They model all system entities within logs as nodes and the interactions between entities as edges, with both nodes and edges bearing attribute information. The nodes within the provenance graph are categorized into subjects and objects based on the direction of data movement. Edges represent the causal relationships between system entities, such as read/write file operations, execute executable file operations, create/clone process operations. By leveraging provenance graphs, security professionals can associate malicious entities with attack behaviors through causal analysis, revealing the complete picture of an attack berrada2020baseline .
2.3 CTI Report and Query Graph
@@
CTI reports openIoc ; stix ; misp encompass comprehensive information related to cyberattacks and attackers, with a particular emphasis on capturing detailed attack procedures, i.e., the intricate sequences of steps and techniques employed in multistage attacks. These reports provide in-depth representations of attack scenarios, potential impacts on target hosts, and the complex chains of causally- linked events that characterize APTs. Security professionals leverage CTI reports to formulate more targeted defense rules for preventing and identifying malicious attack behaviors. In recent years, substantial research altinisik2023provg ; milajerdi2019poirot ; gao2021enabling ; husari2017ttpdrill ; liao2016acing ; lv2024trec has demonstrated the successful application of CTI reports in threat detection and threat hunting. In this paper, we construct a directed graph, called the query graph, from the offensive and defensive knowledge satvat2021extractor (attack entities and their causal relationships) extracted from CTI reports. Similar to provenance graphs, query graphs are directed graphs with attribute information.
2.4 Graph Alignment
@@
Graph alignment refers to the problem of detecting potential cyber intrusion behaviors by establishing an optimal subgraph mapping between a provenance graph ($G_{p}$) representing system activities across the entire system, and a query graph ($G_{q}$) representing attack pattern activities. The provenance graph $G_{p}$ = ($V_{p}$, $E_{p}$) consists of a node set $V_{p}$ representing system entities and events, and an edge set $E_{p}$. Meanwhile, the query graph $G_{q}$ = ($V_{q}$, $E_{q}$) comprises a node set $V_{q}$ representing attack patterns and an edge set $E_{q}$. The goal of graph alignment is to find a subgraph $G_{m}$ in $G_{p}$ that maximizes the matching degree with $G_{q}$:
@@
Here, $M$ is the match function that calculates the matching degree between $G_{q}$ and a subgraph $G^{\prime}$ of $G_{p}$. The best mapping from the attack query graph to the activity graph is obtained by solving this optimization problem, enabling the detection and tracking of cyber intrusion behaviors.
3 RelatedWork
3.1 Provenance Graph-based Threat Hunting
The provenance graph contains causal relationships between system events, and is a data structure that can be effectively utilized for cyber threat hunting. DeepHunter wei2021deephunter utilizes neural tensor networks to judge the subgraph relationship through graph embedding. However, due to the need for multiple-to-multiple traversal comparisons between subgraphs, the efficiency of DeepHunter will decrease as the size of the provenance graph increases. ProvG-Searcher altinisik2023provg establishes a coarse-grained threat hunting by employing a graph representation learning method on the subgraph, with the aim to enhance hunting efficiency. ThreatRaptor gao2021enabling extracts structured threat behaviors from OSCTI and automatically synthesizes query statements to search for malicious activities. Unlike the above studies, Actminer, inspired by POIROT milajerdi2019poirot , tries to align attack graphs extracted from CTI reports with provenance graphs from system logs to find complete attack chain. However, Actminer significantly reduces FPs, FNs, and system overhead through an optimized graph alignment algorithm.
3.2 Graph Matching Algorithms
Graph pattern matching refers to the problem of finding similarities between a small query graph and a large target graph, with widespread applications (e.g., database analysis, search engines, software plagiarism detection, and social networks). Based on whether the matching results are completely consistent, algorithms can be divided into two categories: graph isomorphism matching and graph approximate matching. The main idea of graph isomorphism matching algorithms is to iteratively map nodes from the query graph to the target graph one by one. Ullmann et al. ullmann1976algorithm propose a backtracking-based algorithm that enumerates all subgraphs that satisfy the matching requirements using a depth-first search approach. However, the enumeration range expands as the scale of the graph gradually increases, resulting in low algorithm efficiency. To address this issue, Cordella et al. cordella2004sub propose the VF2 algorithm, which improves algorithm efficiency by incorporating the verification order of query nodes. However, in practical applications, its time complexity is superlinear, and the need for secondary filtering consumes a significant amount of time. Shasha et al. shasha2002algorithmics introduce the GraphGrep algorithm, which achieves fast matching by encoding node semantic information. Several other studies cheng2007fg ; zhao2007graph ; yan2004graph use graph mining techniques to find subgraphs from databases and then employ filtering and optimization strategies to prune incorrect nodes. However, we cannot always assume that the nodes and edges of the query graph can be fully mapped to the target graph due to the diversity and complexity of APT techniques. Graph approximate matching algorithms often rely on heuristic methods to identify important nodes and then gradually expand to neighboring nodes milajerdi2019poirot ; tian2007saga ; tian2008tale ; tong2007fast . Tian et al. tian2007saga use a graph distance model to measure the similarity between graphs. He et al. he2006closure introduce an index-based algorithm to support subgraph queries and similarity queries. Several other works pohly2012hi ; tong2007fast consider the shape and edge attributes of the query graph. However, the aforementioned research overlooks adversarial knowledge. Milajerdi et al. milajerdi2019poirot propose POIROT, which is similar to our work and utilizes node attributes and information flows between nodes for approximate matching. However, it may result in missed detections if an attacker intentionally takes a detour to achieve their goal. Furthermore, if the similarity score exceeds the threshold, POIROT stops hunting, and the obtained attack subgraph may not represent the optimal attack behavior. We develop a new matching technique to address these challenges.
3.3 Incremental Graph Computation
In real-world scenarios, graphs are typically large in scale and frequently updated over time. When graphs are updated, traditional batch processing methods require starting the computation from scratch, which is extremely time-consuming. In contrast to traditional batch processing algorithms, incremental matching only analyzes and matches the updated portion, utilizing previous matching results to maximize the reduction of redundant computations. Fan et al. propose the IncSIMMatch fan2013incremental algorithm, which effectively reduces redundant computations by creating an index for the pattern graph. They then introduce an incremental computation method IncISO fan2017incremental , where only the set of nodes within $d$ hops of the updated node $d$ in the data graph needs to be rematched as the affected region when nodes and edges change in the graph. Subsequently, researchers kao2016distributed ; choudhury2015selectivity ; kankanamge2017graphflow ; idris2017dynamic ; idris2020general ; kim2018turboflux ; min2021symmetric introduce the idea of incremental computation and incorporated query search optimization strategies. For example, Sutanay et al. choudhury2015selectivity construct the pattern graph as a binary tree and decompose it for storage in tree nodes, with data updates performed by searching leaf nodes. However, storing a large number of indexes consumes memory. Sun et al. sun2022depth propose an exploration-based approximate matching technique; however, if the initially selected node is inappropriate, a large number of useless intermediate result values will be generated. Actminer constructs a tree structure to incrementally update the alignment results and introduces a forgetting rate to maintain stable memory overhead.
4 Motivation
4.1 Motivating Example
@@
As illustrated in Figure 1, this example includes two graphs: the top left depicts the attack query graph manually extracted from a CTI report, following the approach outlined in the POIROT.
The attack initiates by leveraging a malicious executable, malicious.exe, to obtain unauthorized system access. It then employs Registry modifications to establish persistence. Once a foothold is secured, the attacker can remotely issue commands and execute them on the system, executing an art.bat_2 file in the temporary folder and facilitating actions such as exfiltrating sensitive data to an external IP, exemplified by the transfer of /etc/passwd containing user account information. This example traces a complete attack chain—from the initial execution of malicious.exe, through Registry-based persistence establishment, to the final remote retrieval of /etc/passwd.
The right side depicts the provenance graph constructed from actual system logs capturing the observed execution behavior. The fragmented nature of attack scenarios, coupled with the constraint of limiting the hop count to existing threat hunting approaches, can lead to imprecise or incomplete results during the threat hunting process. In this paper, we transform the threat hunting problem into finding the attack query graph within the provenance graph.
@@
False Negative. Semantic gaps exist between provenance graphs and attack query graphs due to the complexity of real enterprise environments. The manifestations of the same attack type may differ across systems, and attackers may utilize the same tools in various ways. For instance, entity names in the attack query graph may have varying representations in the underlying logs of different systems. In POIROT, regular expressions are employed to instantiate node names from the attack query graph for hunting searches in the provenance graph. However, if attackers modify their tactics, introducing technical variations, threat hunting systems struggle to detect different mutated attack samples (e.g., over 100 versions of the Carbanak malware were described in CTI reports). Furthermore, attackers can evade security detection through obfuscation, persistence, and evasion techniques. As illustrated, the attack query graph only describes data exfiltration over the network, whereas in the real environment, the attacker creates processes to read files, followed by another process reading those files, and finally transmitting them over the network. Overreliance on the simplistic threat hunting approach based on a predetermined number of hops may inadvertently overlook malicious activities that align with the intrinsic characteristics of attacks, ultimately resulting in detection effort failure. As shown in Figure 1, in a simulated scenario, as described in milajerdi2019poirot Section 5, we set $C_{thr}$ to 3 but find that this limitation resulted in an incomplete capture of the attack chain. Consequently, it was unable to detect the art_bat file. Moreover, with such hop count restrictions, attackers that are aware of the imposed limits could potentially evade hunting more easily across different scenarios, resulting in potential harm. The manual adjustment of the hop limit according to different scenarios poses significant challenges. Similarly, the same problem exists in other path-based detection hassan2019nodoze ; hassan2020tactical ; wang2020you efforts.
@@
False Positive. Within real organizations, extensive legitimate user operations exhibit similarities with attack behaviors in log data. Normal behaviors may be misclassified as malicious if hunting rules are overly broad or incomplete. For example, when a user downloads network files through a browser, the browser collects user data and transmits it to its cloud server, while the downloaded network files may be flagged as “suspicious files” by the system, resembling malicious attack behaviors and triggering false alarms from the hunting system. Additionally, attackers may leverage tools/techniques to deceive hunting systems, which also leads to FPs. As highlighted in Figure 1, a suspicious process Mal.exe exhibits two paths for reading/writing sensitive files. According to the attack query graph, sensitive file access should occur before network transmission, while $path_{1}$ occurs after transmission. Therefore, $path_{2}$ represents the attacker’s actual operations in the environment.
@@
High Overhead and Inefficiency. Government and enterprise organizations typically need to collect data simultaneously from thousands of machines, easily amassing petabyte-scale data volumes. This massive data not only imposes substantial storage overhead but also significantly reduces hunting efficiency. Traditional hunting methods require offline storage and continuous back scanning of system log data, resulting in immense computational overhead for each hunting operation. Referring to Figure 1, assume that all operations before node E23.txt at time 114 have already occurred. When a security analyst attempts to hunt for threats solely based on the data collected after this time, the incremental data segment alone cannot effectively support the reconstruction of the complete attack chain represented by its query graph. When examining the issue holistically, the newly acquired data lacks the necessary evidence to capture the earlier stages of the multistep intrusion. Consequently, subsequent data collection would necessitate rescanning the previously available information, redundantly recomputing the provenance of data prior to a specific time frame. These redundant computations across multiple hunting activities introduce an unsustainable overhead, hindering the efficiency and scalability of the system and potentially allowing malicious activities to persist undetected for extended periods.
5 System Design